### v1

@@ 5.5 M67 (Collinder 204, MWSC 1585, Melotte 94, NGC 2682, OCL 549)
Tarricq et al. (2022) reported halo stars but no tidal tails for M67, while Gao (2020a) identified 1618 likely members, including 85 extra-tidal stars beyond their tidal radius ($r_{t}$) of 16.3 pc, and observed two tidal tails of $\approx$39 pc.
In our analysis, we detected 1200 stars: 1037 cluster members, 163 tail stars, 110 extra-tidal stars (beyond $r_{t}=16.3$ pc), and 73 extra-tidal stars (beyond $r_{J}=20.6$ pc). The leading and trailing tails span $\approx$29 pc and $\approx$31 pc, respectively, with a total tip-to-tip extent of $\approx$102 pc. This extent is slightly shorter than in some earlier reports, likely because previous estimates relied only on 2D sky projections. We account for the cluster’s 3D orientation and proper motions to more conservatively identify tail members. This reduces contamination from foreground or background stars that may align by chance in projection, and avoids overestimating tail lengths based on stars not dynamically connected to the cluster. The orientation of the observed tails along the orbit, with tail stars at an angle, matches the theoretical expectations (Figure 2). Figure 3 reveals that in the cluster core (blue vectors), motion is predominantly oriented perpendicular to the orbit, lacking a clear rotational pattern. Tail vectors (black and red) show a slight anticlockwise motion, indicating weak rotation, with tail stars moving away from the expected orbital path.
Figure 11 (h) presents the luminosity functions (LFs) for the cluster and its tidal tails, with a KS statistic of 0.10 and a p-value of 0.14, indicating no significant difference. A slight excess of low-mass stars in the tails suggests mild mass segregation, but overall, the stellar distributions remain comparable. The binary fraction in the cluster is found to be $0.28\pm 0.02$, while the tails show a slightly higher average value of $0.30$. The leading and trailing tails exhibit binary fractions of $0.33\pm 0.09$ and $0.26\pm 0.08$, respectively.
@@ 6 Cluster Rotation
Our analysis draws inspiration from the residual velocity method outlined in the study, (Hao et al., 2022, 2024), and the methodology proceeds as follows:
@@
To validate our method, we applied it to the Praesepe cluster using data from Hao et al. (2022) and compared the results with their reported values. Our derived azimuthal velocities were $v_{\varphi}=0.1\pm 0.07$ km s-1 within their 1 tidal radius ($r_{t}$) of 10 pc and $0.2\pm 0.07$ km s-1 for the full sample, in close agreement with Hao et al. (2022, 2024). The $v_{\varphi}$ and RMS $v_{\varphi}$ profiles as a function of radius also showed consistency. Using their core radius estimates ($r_{\mathrm{co}}=1.6,\ 2.0,\ 2.6\,\mathrm{pc}$), we computed tidal masses of $M_{t}=(476\pm 120),\ (521\pm 132),\ (581\pm 150)\,M_{\odot}$, respectively. These are also similar to the previously reported values of $M_{t}=(537\pm 146),\ (609\pm 135),\ (\approx 510)\,M_{\odot}$, respectively, supporting the validity of our mass estimation method.
@@ 6.1 Sample of the Stars
We identified cluster and tail stars with reliable radial velocity (RV) values, which in our case are provided primarily for brighter stars with apparent magnitudes of $G<15$ and stars with RV errors smaller than 5 km s-1. In the case of M67, we obtained a sample of 458 stars with reliable RV measurements. The dataset for NGC 2281 was significantly smaller, containing only 123 stars with reliable RVs. Meanwhile, NGC 2354, Alessi 2, and BH 164 were excluded from the analysis, as their reliable RV data had very few stars. Such a small sample size was inadequate for establishing reliable statistical trends or drawing meaningful astrophysical conclusions about the clusters’ kinematics.
@@ 6.2 Coordinate Systems and Rotational Analysis
    The projection of the rotation axis $\overrightarrow{l}$ onto the $Y_{c}$-$Z_{c}$ plane ($l_{yz}$) divides the cluster into two sub-samples. For a rotating cluster, these sub-samples should exhibit opposite signs in their mean residual velocities $v_{x_{c}}$.
@@
    Starting from $\alpha=0^{\circ}$ (aligned with the $Y_{c}$-axis), we incrementally rotate $l_{yz}$ counterclockwise, computing the mean residual velocities $v_{x_{c}}$ for both sub-samples at each angle.
@@
    The optimal angle $\alpha$ is identified where the difference in mean residual velocities between the sub-samples reaches its extremum, indicating the true projection of $\overrightarrow{l}$ onto the $Y_{c}$-$Z_{c}$ plane.
@@
    The process is repeated to determine $\beta$ and $\gamma$, with these angles satisfying the relationship: $\tan\alpha\cdot\tan\gamma=\tan\beta$
@@
Figure 5: The rotational velocity component as a function of the distance $r$ from the cluster center for the cluster + tail stars within 120 pc of M67. Similar plot for NGC 2281 in 12.
@@ 6.3 M67 Rotation
Using ($\alpha,\beta,\gamma$) = ($43.3^{\circ},127.7^{\circ},126.1^{\circ}$), the mean azimuthal velocity $v_{\varphi}$ was found to be $0.84\pm 0.07\,\mathrm{km\,s^{-1}}$ for the entire cluster + tail system, $0.81\pm 0.06\,\mathrm{km\,s^{-1}}$ for the cluster-alone, and $0.89\pm 0.43\,\mathrm{km\,s^{-1}}$ for the tails-alone.
To study the rotational properties of the cluster in depth, we used the Plummer model (Plummer, 1915) to fit the RMS $v_{\varphi}$ versus $r$ curve, which is given by:
$\rho(r)=\frac{3M_{J}}{4\pi r_{\mathrm{co}}^{3}}\left[1+\left(\frac{r}{r_{\mathrm{co}}}\right)^{2}\right]^{-5/2}$ (7)
Here, $M_{J}$ is the total mass enclosed within the Jacobi radius of the cluster, and $r_{\mathrm{co}}$ is the core radius. The corresponding enclosed mass $M(r)$ and circular speed $v_{c}(r)$ as functions of radius are expressed as:
$M(r)=M_{J}\cdot\frac{r^{3}}{\left(r^{2}+r_{\mathrm{co}}^{2}\right)^{3/2}},$ (8)
$v_{c}(r)=\left[\frac{GM_{t}\cdot r^{2}}{\left(r^{2}+r_{\mathrm{co}}^{2}\right)^{3/2}}\right]^{1/2}\quad\Rightarrow\quad v_{c}(r)=v_{0}\cdot\frac{r}{\left(r^{2}+r_{\mathrm{co}}^{2}\right)^{3/4}},$ (9)
where we defined the constant $v_{0}=\sqrt{GM_{J}}$. We fitted the Plummer model to the observed rotation curve of M67 (Figure 6), using literature values for the core radius. Hunt & Reffert (2024) reported $r_{\mathrm{co}}=2.6\,\mathrm{pc}$, while Tarricq et al. (2022) reported $r_{\mathrm{co}}=1.9\,\mathrm{pc}$. For $r_{\mathrm{co}}=2.6\,\mathrm{pc}$, we found a best-fit mass within the Jacobi radius of $M_{J}=(876\pm 144)\,M_{\odot}$, whereas for $r_{\mathrm{co}}=1.9\,\mathrm{pc}$, the best-fit mass within the Jacobi radius was $M_{J}=(832\pm 137)\,M_{\odot}$. Integrated mass of M67 reported by Almeida et al. (2023) is $(1843\pm 368)\ M_{\odot}$ and mass within the Jacobi radius reported by Hunt & Reffert (2024) is $(2779\pm 399)\ M_{\odot}$.
@@ 6.4 NGC 2281 Rotation
The mean residual velocities of the sub-samples divided by the rotation axis showed consistent opposite signs in the cluster + tail and cluster-alone populations, clearly visible only in the $Y_{c}Z_{c}$ and $X_{c}Z_{c}$ planes (Figure 12 a). The $X_{c}Y_{c}$ plane showed minimal indication of rotation, so the angle $\gamma$ was unreliable in this case. The tails-alone population lacked a clear rotational signature and appeared to be dominated by noise. Within one Jacobi radius ($r_{J}$) of NGC 2281, the RMS $v_{\varphi}$ followed Newtonian expectations but showed significant dispersion (Figure 12 b), which decreased beyond that Jacobi limit.
@@
We estimated the mass within the Jacobi radius using the Plummer model fit to the RMS $v_{\varphi}$ vs. radius curve (Figure 12 c), considering two core radii: $r_{\mathrm{co}}=2.6\,\mathrm{pc}$ (Hunt & Reffert, 2024) and $r_{\mathrm{co}}=1.5\,\mathrm{pc}$ (Tarricq et al., 2022). The best-fit masses were $M_{J}=(803\pm 296)\,M_{\odot}$ for $r_{\mathrm{co}}=2.6\,\mathrm{pc}$ and $(741\pm 268)\,M_{\odot}$ for $r_{\mathrm{co}}=1.5\,\mathrm{pc}$. Integrated mass of NGC 2281 reported by Almeida et al. (2023) is $(681\pm 136)\ M_{\odot}$ and mass within the Jacobi radius reported by Hunt & Reffert (2024) is $(885\pm 112)\ M_{\odot}$.
@@ 7 Discussion
We selected five open clusters with elongated morphology, as suggested by (Bhattacharya et al., 2022; Tarricq et al., 2022; Hunt & Reffert, 2023, 2024; Gao, 2020a), and used DBSCAN to identify their member stars. To detect tidal tails, we first examined specific properties indicative of tail structures, including RV and CMD. We then applied an unsupervised machine learning approach, incorporating PCA to determine the elongation axis, followed by DBSCAN to identify the tails using Gaia DR3 data. Our analysis revealed that all five clusters exhibit tail stars extending beyond their Jacobi radii (Figure 2). Additionally, we accounted for projection effects to refine the spatial distribution of the tails (Subsection 3.4) and their proper motion vectors (Subsection 4.1). Figures 7 to 11 show detailed plots for each cluster, including their spatial distributions, color-magnitude diagrams, vector point diagrams, and luminosity functions for both cluster members and tidal tails.
There is a notable increase in the number of stars identified in tidal tails across the clusters compared to what is documented in literature (Bhattacharya et al., 2022; Tarricq et al., 2022; Hunt & Reffert, 2023, 2024; Gao, 2020a) with higher star count in four clusters (BH 164, Alessi 2, NGC 2281, and NGC 2354) and longer tidal tails in three (Alessi 2, NGC 2281, and NGC 2354). The improvements vary from case to case: star counts increased in our study by as few as 29 stars (Alessi 2) to as many as 131 (NGC 2281) as compared to the literature, while the tip-to-tip extent in our study increased by 8 pc (NGC 2354) up to 43 pc (Alessi 2) as compared to the literature. In M67, although we identified 1200 stars in total, the observed span of 102 pc was slightly shorter than that reported by Gao (2020a).
@@
This difference may be due to our consideration of projection effects, which refines the estimation of tail lengths in the observed plane. Among the clusters studied, M67 exhibited the longest tidal tail span, followed by NGC 2281, Alessi 2, NGC 2354, and BH 164. The youngest cluster (BH 164, 65 Myr) showed the shortest tidal tail span, while the oldest (M67, 4 Gyr) exhibited the longest. According to N-body simulations (Jerabkova et al., 2021), older clusters have longer tails; however, we did not detect complete tails ($\leq 40$ pc). Hence, the detected trend is a combination of the actual tail span, the different populations, and the incompleteness of the detection.
@@
In addition, the clustering algorithms we used in this work select stars with similar CMD location and motion as tail candidates. However, such criteria can also be satisfied by non-tail stars, leading to contamination of the sample, as seen in the clump we removed in NGC 2281 and in the unusually extended leading tail of NGC 2354 compared to simulations. Similar unexpected structures are also seen in other tail catalogues, for example, in the Pleiades (Risbud et al. 2025), which correspond to comoving stars but might not be real tails. More analysis is required to understand the completeness and purity of the tail candidates.
Proper motion vectors provided information about the motion of both the cluster and tail stars (Section 5), shedding light on the overall dynamical state. These patterns indicated uniform alignment toward a specific direction, random disordered motion, or potential rotation. The luminosity functions for all the clusters revealed that their tails notably lacked high-mass stars, which is expected behavior.
@@
When comparing the average BF of tidal tails to cluster cores, we found that in 3 out of 5 clusters, the tails exhibited higher binary fractions. Specifically, Alessi 2 showed the most pronounced difference, with tails averaging BF = 0.47 compared to the cluster’s BF = 0.22. NGC 2281 (tails: average BF = 0.28 vs. cluster: BF = 0.24) and M67 (tails: average BF = 0.30 vs. cluster: BF = 0.28) also followed this trend. BH 164 showed equal binary fractions (BF = 0.26) in the tails and the cluster, while NGC 2354 was the only cluster in which the tails showed a lower average binary fraction (BF = 0.27) than the cluster (BF = 0.34). This generally higher binary fraction in tidal tails may result from dynamical interactions preferentially ejecting binary systems from cluster cores. All tidal tails demonstrated a clear deficiency of high-mass stars compared to their respective cluster centers. This deficiency aligns with theoretical predictions of mass segregation, where higher-mass stars sink toward cluster centers during dynamical evolution, while lower-mass stars and binaries are more susceptible to populating the outer regions and tidal tails.
@@
Based on simulations, the escaping stars from the tidal tails have preferred velocities where the leading tail stars move towards the Galactic center and the trailing tail stars move away from the Galactic center. This can be observed as an apparent rotation in the sky and the galactic plane. We examined the clusters and their tidal tails using the residual velocity method to investigate their motion dynamics and 3D rotational behavior. We detected clear signatures of rotation in M67 and NGC 2281. However, BH 164, Alessi 2, and NGC 2354 were excluded from the rotation analysis due to insufficient reliable radial velocity data. M67 demonstrated rotation not only in its central cluster but also extending into its tidal tails, with a mean ($v_{\varphi}$) of $0.84\pm 0.07$ km s-1 for the cluster and tail system. This rotational signature of M67, along with the expected sky positions, confirms that the extended structure around M67 is indeed its tidal tail. In contrast, NGC 2281 showed rotation primarily in its cluster region with $v_{\varphi}=0.53\pm 0.13$ km s-1, while its tails exhibited minimal contribution to the overall rotation.

### v3

@@ 5.5 M67 (Collinder 204, MWSC 1585, Melotte 94, NGC 2682, OCL 549)
Tarricq et al. (2022) reported halo stars but no tidal tails for M67, while Gao (2020a) identified 1618 likely members, including 85 extra-tidal stars beyond their tidal radius of 16.3 pc, and observed two tidal tails of $\approx$39 pc. In our analysis, we detected 1200 stars: 1037 cluster members, 163 tail stars, 110 extra-tidal stars (beyond the tidal radius of 16.3 pc), and 73 extra-tidal stars (beyond $r_{J}=20.6$ pc). The leading and trailing tails span $\approx$29 pc and $\approx$31 pc, respectively, with a total tip-to-tip extent of $\approx$102 pc. This extent is slightly shorter than in some earlier reports, likely because previous estimates relied only on 2D sky projections. We account for the cluster’s 3D orientation and proper motions to more conservatively identify tail members. This reduces contamination from foreground or background stars that may align by chance in projection, and avoids overestimating tail lengths based on stars not dynamically connected to the cluster. The orientation of the observed tails along the orbit, with tail stars at an angle, matches the theoretical expectations (Fig. 2). Figure 3 reveals that in the cluster core (blue vectors), motion is predominantly oriented perpendicular to the orbit, lacking a clear rotational pattern. Tail vectors (black and red) show a slight counterclockwise motion, indicating weak rotation, with tail stars moving away from the expected orbital path.
Figure 11 (h) presents the LFs for the cluster and its tidal tails, with a KS statistic of 0.10 and a p-value of 0.14, indicating no significant difference. A slight excess of low-mass stars in the tails suggests mild mass segregation, but overall, the stellar distributions remain comparable. The binary fraction in the cluster is found to be $0.28\pm 0.02$, while the tails show a slightly higher average value of $0.30$. The leading and trailing tails exhibit the binary fractions of $0.33\pm 0.09$ and $0.26\pm 0.08$, respectively.
@@ 6 Cluster rotation
Our analysis draws inspiration from the residual velocity method outlined in the study by Hao et al. (2022, 2024), and the methodology proceeds as follows:
@@
To validate our method, we applied it to the Praesepe cluster using data from Hao et al. (2022) and compared the results with their reported values. Our derived mean azimuthal velocity ($v_{\varphi}$) is $0.1\pm 0.07$ km s-1 within the tidal radius of 10 pc and $0.2\pm 0.07$ km s-1 for the full sample, in close agreement with Hao et al. (2022, 2024). The $v_{\varphi}$ and root mean squared (RMS) $v_{\varphi}$ profiles as a function of radius also showed consistency. Using their core radius estimates ($r_{\mathrm{co}}=1.6,\ 2.0,\ 2.6\,\mathrm{pc}$), we computed tidal masses of $M_{t}=(476\pm 120),\ (521\pm 132),\ (581\pm 150)\,M_{\odot}$, respectively. These are also similar to the previously reported values of $M_{t}=(537\pm 146),\ (609\pm 135),\ (\approx 510)\,M_{\odot}$, respectively, supporting the validity of our mass estimation method.
@@ 6.1 Sample of the stars
We identified cluster and tail stars with reliable RV values, which in our case are provided primarily for brighter stars with apparent magnitudes of $G<15$ and stars with RV errors smaller than 5 km s-1. In the case of M67, we obtained a sample of 458 stars with reliable RV measurements. The dataset for NGC 2281 was significantly smaller, containing only 123 stars with reliable RVs. Meanwhile, NGC 2354, Alessi 2, and BH 164 were excluded from the analysis, as their reliable RV data had very few stars. Such a small sample size was inadequate for establishing reliable statistical trends or drawing meaningful astrophysical conclusions about the clusters’ kinematics.
@@ 6.2 Coordinate systems and rotational analysis
    The projection of the rotation axis $\overrightarrow{l}$ onto the $Y_{c}$-$Z_{c}$ plane ($l_{yz}$) divides the cluster into two subsamples. For a rotating cluster, these subsamples should exhibit opposite signs in their mean residual velocities $v_{x_{c}}$.
@@
    Starting from $\alpha=0^{\circ}$ (aligned with the $Y_{c}$-axis), we incrementally rotate $l_{yz}$ counterclockwise, computing the mean residual velocities $v_{x_{c}}$ for both subsamples at each angle.
@@
    The optimal angle $\alpha$ is identified where the difference in mean residual velocities between the subsamples reaches its extremum, indicating the true projection of $\overrightarrow{l}$ onto the $Y_{c}$-$Z_{c}$ plane.
@@
    The process is repeated to determine $\beta$ and $\gamma$, with these angles satisfying the relationship: $\tan\alpha\tan\gamma=\tan\beta$.
@@
Figure 5: Rotational velocity component as a function of the distance $r$ from the cluster center for the cluster + tail stars within 120 pc of M67. Similar plot for NGC 2281 in Fig. 12.
@@ 6.3 M67 rotation
To study the rotational properties of the cluster in depth, we used the Plummer model (Plummer, 1915) to fit the RMS $v_{\varphi}$ versus distance $r$ curve, which is as follows:
$\rho(r)=\frac{3M_{J}}{4\pi r_{\mathrm{co}}^{3}}\left[1+\left(\frac{r}{r_{\mathrm{co}}}\right)^{2}\right]^{-5/2}.$ (6)
Here, $M_{J}$ is the total mass enclosed within the $r_{J}$ of the cluster, and $r_{\mathrm{co}}$ is the core radius. The corresponding enclosed mass $M(r)$ and circular speed $v_{c}(r)$ as functions of radius are expressed as
$M(r)=M_{J}\frac{r^{3}}{\left(r^{2}+r_{\mathrm{co}}^{2}\right)^{3/2}},$ (7)
$v_{c}(r)=\left[\frac{GM_{t}\ r^{2}}{\left(r^{2}+r_{\mathrm{co}}^{2}\right)^{3/2}}\right]^{1/2}\quad\Rightarrow\quad v_{c}(r)=v_{0}\frac{r}{\left(r^{2}+r_{\mathrm{co}}^{2}\right)^{3/4}},$ (8)
where we defined the constant $v_{0}=\sqrt{GM_{J}}$. We fit the Plummer model to the observed rotation curve of M67 (Fig. 6), using literature values for the core radius. Hunt & Reffert (2024) reported $r_{\mathrm{co}}=2.6\,\mathrm{pc}$, while Tarricq et al. (2022) reported $r_{\mathrm{co}}=1.9\,\mathrm{pc}$. For $r_{\mathrm{co}}=2.6\,\mathrm{pc}$, we found a best-fit mass within the $r_{J}$ of $M_{J}=(876\pm 144)\,M_{\odot}$, whereas for $r_{\mathrm{co}}=1.9\,\mathrm{pc}$, the best-fit mass within the $r_{J}$ was $M_{J}=(832\pm 137)\,M_{\odot}$. Integrated mass of M67 reported by Almeida et al. (2023) is $(1843\pm 368)\ M_{\odot}$ and mass within the $r_{J}$ reported by Hunt & Reffert (2024) is $(2779\pm 399)\ M_{\odot}$.
@@ 6.4 NGC 2281 rotation
[…]
We estimated the mass within the $r_{J}$ using the Plummer model fit to the RMS $v_{\varphi}$ vs. distance r curve (Fig. 12), considering two core radii: $r_{\mathrm{co}}=2.6\,\mathrm{pc}$ (Hunt & Reffert, 2024) and $r_{\mathrm{co}}=1.5\,\mathrm{pc}$ (Tarricq et al., 2022). The best-fit masses were $M_{J}=(803\pm 296)\,M_{\odot}$ for $r_{\mathrm{co}}=2.6\,\mathrm{pc}$ and $(741\pm 268)\,M_{\odot}$ for $r_{\mathrm{co}}=1.5\,\mathrm{pc}$. Integrated mass of NGC 2281 reported by Almeida et al. (2023) is $(681\pm 136)\ M_{\odot}$ and mass within the $r_{J}$ reported by Hunt & Reffert (2024) is $(885\pm 112)\ M_{\odot}$.
@@ 7 Discussion
We selected five open clusters with elongated morphology, as suggested by (Bhattacharya et al., 2022; Tarricq et al., 2022; Hunt & Reffert, 2023, 2024; Gao, 2020a), and used DBSCAN to identify their member stars. To detect tidal tails, we first examined specific properties indicative of tail structures, including RV and the CMD. We then applied an unsupervised machine learning approach, incorporating PCA to determine the elongation axis, followed by DBSCAN to identify the tails using Gaia DR3 data. Our analysis revealed that all five clusters exhibit tail stars extending beyond their Jacobi limit (Fig. 2). We corrected the proper motion vectors for projection effects (Subsect. 4.1) and refined the spatial distribution of the tails by accounting for the spherical geometry of the sky (Subsect. 3.4). Figures 7 to 11 show detailed plots for each cluster, including their spatial distributions, CMDs, vector point diagrams, and LFs for both cluster members and tidal tails.
@@
There is a notable increase in the number of stars identified in tidal tails across the clusters compared to what is documented in literature (Bhattacharya et al., 2022; Tarricq et al., 2022; Hunt & Reffert, 2023, 2024; Gao, 2020a) with higher star count in four clusters (BH 164, Alessi 2, NGC 2281, and NGC 2354) and longer tidal tails in three (Alessi 2, NGC 2281, and NGC 2354). The improvements vary from case to case: star counts increased in our study by as few as 29 stars (Alessi 2) to as many as 131 (NGC 2281), as compared to the literature, while the tip-to-tip extent in our study increased by 8 pc (NGC 2354) up to 43 pc (Alessi 2) as compared to the literature. In M67, although we identified 1200 stars in total, the observed span of 102 pc was slightly shorter than that reported by Gao (2020a). This difference may be due to our consideration of projection effects, which refines the estimation of tail lengths in the observed plane. Among the clusters studied, M67 exhibited the longest tidal tail span, followed by NGC 2281, Alessi 2, NGC 2354, and BH 164. The youngest cluster (BH 164, 65 Myr) showed the shortest tidal tail span, while the oldest (M67, 4 Gyr) exhibited the longest. According to N-body simulations (Jerabkova et al., 2021), older clusters have longer tails; however, we did not detect complete tails ($\leq 40$ pc). Hence, the detected trend is a combination of the actual tail span, the different populations, and the incompleteness of the detection.
@@
In addition, clustering algorithms and statistical methods were used in this work to select stars with similar CMD locations and motions as tail candidates. However, such criteria can also be satisfied by field stars, leading to contamination of the sample, as seen in the clump we removed in NGC 2281 and in the unusually extended leading tail of NGC 2354 compared to simulations. Similar unexpected structures are also seen in other tail catalogs, for example, in the Pleiades (Risbud et al. 2025), which correspond to comoving stars but might not be real tails. More analysis is required to understand the completeness and purity of the tail candidates.
@@
Proper motion vectors provided information about the motion of both the cluster and tail stars (Sect. 5), shedding light on the overall dynamical state. These patterns indicated uniform alignment toward a specific direction, random disordered motion, or potential rotation. The LFs for all the clusters revealed that their tails notably lacked high-mass stars, which is expected behavior. When comparing the average binary fraction of tidal tails to that of cluster cores, we found that in three out of five clusters, the tails contained a higher proportion of binaries. Alessi 2 showed the most pronounced difference, with the tails averaging a binary fraction of 0.47 compared to 0.22 in the cluster core. NGC 2281 (tails: 0.28 vs. cluster: 0.24) and M67 (tails: 0.30 vs. cluster: 0.28) also followed this trend. BH 164 showed equal binary fractions (0.26) in both the tails and the cluster, while NGC 2354 was the only case where the tails had a lower average binary fraction (0.27) than the cluster (0.34). This generally higher binary fraction in tidal tails may result from dynamical interactions preferentially ejecting binary systems from cluster cores. All tidal tails demonstrated a clear deficiency of high-mass stars compared to their respective cluster centers. This deficiency aligns with theoretical predictions of mass segregation, where higher-mass stars sink toward cluster centers during dynamical evolution, while lower-mass stars and binaries are more susceptible to populating the outer regions and tidal tails.
@@
Based on simulations, the escaping stars from the tidal tails have preferred velocities where the leading tail stars move toward the Galactic center and the trailing tail stars move away from the Galactic center. This can be observed as an apparent rotation in the sky and the Galactic plane. We examined the clusters and their tidal tails using the residual velocity method to investigate their motion dynamics and 3D rotational behavior. We detected clear signatures of rotation in M67 and NGC 2281. However, BH 164, Alessi 2, and NGC 2354 were excluded from the rotation analysis due to insufficient reliable RV data. M67 demonstrated rotation not only in its central cluster but also extending into its tidal tails, with a mean $v_{\varphi}$ of $0.84\pm 0.07$ km s-1 for the cluster and tail system. This rotational signature of M67, along with the expected sky positions, confirms that the extended structure around M67 is indeed its tidal tail. In contrast, NGC 2281 showed rotation primarily in its cluster region with $v_{\varphi}=0.53\pm 0.13$ km s-1, while its tails exhibited minimal contribution to the overall rotation.